• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 18:28:48 on July26,2003

    Entry number 107121

    This entry is a followup to: 107096

    Followups:

    keyword=L-arm helicity restored (probably)

    It appears that the 2 signals "Gate" and "Helicity" were swapped at the patch panel.
    I have no confirmation about what is the correct patch panel cabling, and we should
    ask the person who was working there Friday -- or else do an access and double
    check the cabling -- but this makes sense because: a) the ordering of cables at the
    patch is now the same on L-arm and R-arm. I did not touch R-arm. b) The scalers
    make sense now; equal data in plus and minus helicity, and the helicity itself is
    counted in 32nd channel and makes sense; and c) Helicity flag shows up now in
    ROC11 I/O register. Unfortunately it does NOT show up in the ADC, but I guess
    that is a different problem (probably some ADC chaos between experiments, the
    ADC flag was for sure working in May).

    So, modulo the above mentioned concerns, the L-arm helicity is probably ok now.